# Learning and composing of classical music using restricted Boltzmann machines

### v1

@@ Abstract
Recently, software has been developed that uses machine learning to mimic the style of a particular composer, such as J. S. Bach. However, since such software often adopts machine learning models with complex structures, it is difficult to analyze how the software understands the characteristics of the composer’s music. In this study, we adopted J. S. Bach’s music for training of a restricted Boltzmann machine (RBM). Since the structure of RBMs is simple, it allows us to investigate the internal states after learning. We found that the learned RBM is able to compose music.
@@ 1 Introduction
Recent progress in machine learning has spurred its integration into various domains of physics. In particular, machine-learning-based methodologies are increasingly being investigated as powerful tools for data analysis, modeling, and simulation in contemporary physics research [1, 2, 3, 4, 5, 6, 7]. In studies integrating physics and machine learning, there have also been attempts to understand the internal mechanisms of machine learning models from a physical perspective [3, 4, 5]. Among the various machine learning models, the Boltzmann machine is particularly notable for its deep relationship with physics. [8]. The Boltzmann machine is an associative memory model that stores input data using spin interactions and local biases as parameters, and its operating principle is based on statistical mechanics. A restricted Boltzmann machine (RBM) is a simplified variant of the Boltzmann machine in which the network is constrained to two layers, making the learning process more tractable [9, 10]. Applied studies of RBMs include examples such as image generation [11] and video anomaly detection [12]. RBMs have also been employed in physics research, for instance, in studies of quantum systems [13, 6].
The simplicity of the RBM architecture enables us to analyze its internal mechanisms from a statistical physics perspective. This analytical accessibility, which distinguishes RBMs from more complex models, represents a key advantage in the development of explainable artificial intelligence (XAI) systems [14], whose importance is increasing as AI becomes more widely deployed in society. Modern deep learning models often possess highly complex architectures, making it challenging for humans to comprehend the meaning and function of the parameters learned during training. As a result, AI systems are often operated as black boxes, with the reasoning behind their outputs remaining inaccessible to humans—a situation that has been regarded as problematic from the perspective of safety and trust [15]. To address this issue, research has been conducted to analyze trained models and make the basis of their outputs understandable to humans. Such analyses have also been carried out in physics research. One illustrative example of such analysis is the discovery that, when a neural network is trained to learn the temperature and spin configurations of the Ising model, its intermediate layers encode the order parameter of the system. This observation implies that the network spontaneously discovers the physical concept of an order parameter solely from the training data, without explicit supervision [4].
The applications of machine learning have progressed not only in the scientific domain but also in the field of arts, with examples such as style transfer of images by AI trained on the styles of specific painters [16, 17, 18], and automatic music composition by AI [19, 20, 21, 22, 23]. In studies on machine learning-based music composition, a variety of approaches have been explored, such as generating chorales in the style of J. S. Bach [20] and constructing large-scale language models capable of composing music [23], with trained models successfully producing original compositions. Some automatic composition systems have already reached commercial-level deployment [22], indicating that machine learning-based composition has attained a practical stage. In addition to composition, machine learning has been applied to automatic harmonization [21] and automatic accompaniment generation [24], expanding the possibilities for AI-assisted music production in various forms. Despite the large number of successful examples in composition, harmonization, and accompaniment, relatively few studies have analyzed the internal workings of trained music generation models [25], and much remains unknown about how these models internally represent and understand music. As interactive collaboration between humans and artificial intelligence in musical activities becomes increasingly viable, AI systems need to exhibit a level of musical understanding comparable to that of human musicians. Accordingly, elucidating how AI models perceive and process music is a matter of significant importance. A deeper understanding of the internal representations acquired during training may not only enhance the interpretability of these models but also enable their use as analytical tools that offer novel insights into musical structure. Nevertheless, most machine learning models employed for music generation rely on complex deep learning architectures, which pose substantial challenges to the interpretability of their internal mechanisms.
In this study, we demonstrate that music composition is feasible using a restricted Boltzmann machine (RBM), a model with a simple structure that permits relatively straightforward analysis of its internal mechanisms after training. Our objective is to understand, from the perspective of statistical physics, how an RBM that performs a task not typically associated with physics, specifically music composition, internally represents musical data. Although previous work, such as that by Boulanger et al. [19], has applied RBMs to music composition, few studies have explored the use of RBMs in isolation for this purpose. In the present work, we propose a method in which an RBM is trained on piano-roll representations of keyboard music. We then analyze the trained model to examine how it processes and encodes musical information, enabling it to generate compositions.
2 Method
@@ 2.1 RBM
In this study, we adopted an RBM as the model for music composition. An RBM is a Boltzmann Machine with a constrained network structure. A Boltzmann machine is a network connecting many units, each of which has spin degrees of freedom. Each edge has a weight, and each unit has a bias, which defines the energy of the network [26, 8]. The Boltzmann machine is a type of recurrent neural network that learns patterns in input data and can then generate them stochastically. While Boltzmann machines were theoretically interesting, the learning cost increases exponentially as the number of units increases. To address the problem of the learning cost, a restricted Boltzmann machine was proposed [8]. An RBM consists of two types of layers: a visible layer and a hidden layer. It is subject to the constraint that there are no connections between units within the same layer and this constraint allows us to adopt the efficient learning procedure. Each unit has a spin degree of freedom, which can be either Gaussian-type or Bernoulli-type. A Gaussian-type unit can take continuous values ranging from minus infinity to plus infinity, whereas a Bernoulli-type unit is binary and takes only the values 0 or 1. In this study, we adopt a Bernoulli-Bernoulli RBM, in which both the visible and hidden units are of the Bernoulli type[27].
@@
where, $\theta=\{w_{ij},b_{i},c_{i}\}$ are the model parameters [28]. The weights $\boldsymbol{W}=w_{ij}$ represents the interaction between the visible unit $v_{i}$ and the hidden unit $h_{j}$. The parameters $\boldsymbol{b}=b_{i}$ and $\boldsymbol{c}=c_{i}$ represent the biases of the visible and hidden layers, respectively. Given the model parameter $\theta$, the probability of the visible layer being in the state $\boldsymbol{v}$ is given by,
$p(\boldsymbol{v}|\theta)=\frac{\exp\{-E(\boldsymbol{v}|\theta)\}}{Z(\theta)},$ (2)
@@
$Z(\theta)=\sum_{\boldsymbol{v}}\sum_{\boldsymbol{h}}\exp\left(-E(\boldsymbol{v},\boldsymbol{h})|\theta\right).$ (3)
@@
In general Boltzmann machines, computing the gradient of this cost function is intractable, while in the case of RBMs, it can be efficiently carried out using the Contrastive Divergence (CD) method [10]. In this study, we adopt the CD method as the optimization technique. We implemented the RBM model using Python, and by utilizing CuPy, we were able to accelerate the computations through GPGPU processing [29]. The RBM code developed in this study is available on GitHub [30].
@@ 2.2 Dataset
For the training data, we adopted compositions by J. S. Bach. A total of 58 MIDI files were obtained from the Mutopia Project [31], and each file was converted into a black-and-white image representation known as a piano roll. A piano roll is a two-dimensional representation of musical information, where the horizontal axis corresponds to time and the vertical axis corresponds to pitch. Notes are depicted as horizontal bars, with their positions and lengths indicating the timing and duration of each note, respectively. Each pixel value in the piano roll image is either 0 or 1, corresponding to the binary visible units of a Bernoulli-type RBM.
@@
The image size was fixed at $72\times 192$ pixels. The vertical dimension of 72 pixels corresponds to the pitch range from C1 to B6, where C1 denotes the C note in the first octave of a standard 88-key piano (i.e., the lowest C key), and B6 denotes the B note in the sixth octave, one semitone below the highest C (C8). The horizontal dimension of 192 pixels represents time, with 24 pixels corresponding to the duration of one quarter note. This resolution was chosen so that the horizontal pixel count would be divisible by 3, enabling the representation of triplet notes.
@@ 2.3 Music Composition
6: By repeating steps 2 through 5 $T$ times, a binary vector is obtained in which exactly NN elements are set to 1.
@@ 3.3 Music Composition
An example of two-measure music generation using Algorithm 1 is shown in Fig. 7. The figure shows the visible states $\mathbf{v}_{t}$at sampling steps $t=50,100,250,500,750,$ and $1000$. All images exhibit the structure of piano rolls and clearly form both chords and melodies. The time evolutions of the energy of the RBM during image generation is shown in Fig. 7. The energy decreases monotonically up to approximately 500 sampling steps, after which it begins to increase. This suggests that the RBM assigns higher energy when the number of active pixels (notes) is either too small or too large, implying the existence of an optimal number of notes that minimizes the energy.
An example of an eight-measure composition generated using Algorithm 2 is shown in Fig. 8. This image also exhibited a piano roll structure, similar to the two-measure images shown in Fig. 6. Upon listening to the music corresponding to this piano roll, we observe that although some unnatural aspects remain, the result forms a coherent piece. This indicates that even a simple RBM can successfully generate music. The audio of this piece, as well as other compositions generated by the RBM trained in this study, can be found online [32].
@@
Figure 8: Piano roll representation of an 8-measure melody composed by the RBM.
3.4 Analysis of the Hidden Layer Using t-SNE
We confirmed that the RBM trained on piano rolls is capable of music composition. Next, to investigate how the trained RBM internally represents musical structure, particularly whether it exhibits human-like understanding, we examined its response to transposed musical inputs. An RBM compresses the information provided to the visible layer and stores it in the hidden layer, from which the visible state can be reconstructed. This suggests that the hidden layer retains feature representations of the input. Therefore, we analyzed how the hidden layer representation changes when a musical piece is transposed.

### v3

@@ Abstract
We investigate how machine learning models acquire the ability to compose music and how musical information is internally represented within such models. We develop a composition algorithm based on a restricted Boltzmann machine (RBM), a simple generative model capable of producing musical pieces of arbitrary length. We convert musical scores into piano-roll image representations and train the RBM in an unsupervised manner. We confirm that the trained RBM can generate new musical pieces; however, by analyzing the model’s responses and internal structure, we find that the learned information is not stored in a form directly interpretable by humans. This study contributes to a better understanding of how machine learning models capable of music composition may internally represent musical structure and highlights issues related to the interpretability of generative models in creative tasks.
articletype: ARTICLE TEMPLATE
@@ 1 Introduction
Recent advances in machine learning have greatly accelerated the development of AI-based music composition. Many contemporary composition systems are built upon deep learning models (Eck and Schmidhuber, 2002; Choi et al., 2016; Mogren, 2016; Hadjeres et al., 2017; Dong et al., 2018), and in recent years, large-scale generative models inspired by language modeling have demonstrated remarkable capabilities in producing music (Hawthorne et al., 2018; Dhariwal et al., 2020; Agostinelli et al., 2023; Yuan et al., 2024). However, such models are generally complex, and it is not easy for humans to understand which musical features they have learned. As a result, the process of music generation tends to become a black box, making it difficult to interpret the underlying generative principles or to explain the model’s internal behavior (Briot and Pachet, 2017; Wang et al., 2020; Bryan-Kinns et al., 2023). In addition, it has been pointed out that the practical usefulness and influence of such systems on real-world music practitioners are rarely examined in a systematic manner (Sturm et al., 2018). Long before the advent of modern deep learning approaches, statistical and information-theoretic methods were already explored as a means of modeling musical style and structure (Dubnov et al., 2003). These early studies demonstrated that musical regularities can be captured from symbolic data without relying on explicit music-theoretical rules, and they laid important groundwork for subsequent research on machine learning-based music modeling.
From the perspective of explainable AI (XAI), it is important that humans can understand the internal representations upon which a generative model bases its outputs (Bryan-Kinns et al., 2023), as such interpretability forms the foundation of trust between the model and human musicians. In particular, in tasks such as musical analysis, examining the features learned by the model can help us discover latent stylistic characteristics and regularities inherent in musical data. Thus, a generative model can serve not only as a system that automatically produces music, but also as an analytical tool that reveals the underlying structure of musical data. Despite this broad recognition, systematic studies that directly analyze the internal mechanisms of models after they have acquired the ability to compose music remain limited. The present work is designed to address this gap.
In this study, we deliberately focus on generative models with transparent and straightforward structures, whose internal states are more amenable to systematic analysis. Rather than aiming to develop a highly optimized composition system, we seek to design an algorithm that enables music generation with a minimal, interpretable model architecture. The central objective is to investigate how such a simple model can acquire the ability to generate music, and to analyze its responses and internal representations once this ability has emerged. By prioritizing interpretability over sheer generative performance, this study adopts a constructive perspective aligned with the goals of explainable AI, treating the generative model not only as a creative system but also as a tool for probing the internal mechanisms of machine learning-based musical representation.
To this end, we adopt the Restricted Boltzmann Machine (RBM) (Smolensky et al., 1986) as our modeling framework, as its simple, constrained architecture is well suited to our goal of analyzing internal representations transparently. An RBM is a probabilistic generative model composed of a visible layer and a hidden layer, and it learns the probability distribution of training data (Ackley et al., 1985; Zhang et al., 2018). In contrast to general Boltzmann machines, the RBM imposes restrictions on network connectivity, resulting in a simpler learning algorithm and a set of internal parameters that are more directly amenable to systematic analysis. Although a standard RBM cannot explicitly model the temporal structure of input sequences, and extensions such as the Temporal RBM (TRBM) (Sutskever and Hinton, 2007) and Conditional RBM (CRBM) (Taylor et al., 2006) have been proposed to address this limitation, we deliberately refrain from introducing such temporal mechanisms. By doing so, we prioritize interpretability and generate music strictly within the standard RBM framework.
Although several studies have applied RBMs to music modeling (Boulanger-Lewandowski et al., 2012; Lyu et al., 2015), many of them employ RBMs as components of deeper architectures, such as recurrent or temporally extended models, rather than examining how well a standard RBM alone can perform musical generation or what kinds of internal representations it acquires. Consequently, there have been few investigations into the extent to which a standalone RBM can generate music and what structural properties its latent representations possess. Recent work has demonstrated that RBMs can successfully generate classical music by training on real musical datasets (Carbone et al., 2025). However, while these studies show the generative capability of RBMs, the mechanisms by which the trained models internally represent and understand musical structure remain unexplored.
In this study, we train a restricted Boltzmann machine (RBM) on piano-roll images derived from keyboard works by J. S. Bach and conduct a multifaceted analysis of its behavior. Specifically, (i) we evaluate how accurately the trained RBM reconstructs the training piano-rolls and analyze the energy values it assigns to unseen piano-rolls and to non-musical images, such as MNIST digits, thereby assessing its ability to distinguish musical from non-musical data. Furthermore, (ii) we generate new piano-roll samples from the trained RBM using the proposed generation algorithm and examine the extent to which it can produce musically coherent structures in sequences of two measures or longer. In addition, focusing on the internal representations of the RBM, (iii) we input one-hot vectors to individual hidden units and analyze the expected visible-layer patterns to investigate the types of musical patterns encoded by each hidden unit.
Through these analyses, we aim to clarify what kinds of musical regularities a standard RBM, here used as a minimal generative model, can learn from piano-roll images of J. S. Bach’s compositions, and whether its internal representations correspond to concepts familiar to human music theory. The results demonstrate that RBMs are capable of musical generation and further suggest that the latent space learned by an RBM may function as a data-driven analytical representation that is not necessarily aligned with conventional music-theoretical frameworks. Taken together, these findings contribute to a foundational understanding of explainable generative models for music.
The remainder of this paper is organized as follows. Section 2 describes the proposed methods. Section 3 presents the experimental results. Finally, Section 4 provides a summary of the findings and discusses their implications.
2 Methods
@@ 2.1 RBM
In this study, we adopted an RBM as the model for music composition. An RBM is a Boltzmann Machine with a constrained network structure. A Boltzmann machine is a network connecting many units, each of which has spin degrees of freedom. Each edge has a weight, and each unit has a bias, which defines the energy of the network (Hinton and Sejnowski, 1983; Ackley et al., 1985). The Boltzmann machine is a type of recurrent neural network that learns patterns in input data and can then generate them stochastically. While Boltzmann machines were theoretically interesting, the learning cost increases exponentially as the number of units increases. To address the problem of the learning cost, a restricted Boltzmann machine was proposed (Ackley et al., 1985). An RBM consists of two types of layers, a visible layer and a hidden layer. It is subject to the constraint that there are no connections between units within the same layer and this constraint allows us to adopt the efficient learning procedure. Each unit has a spin degree of freedom, which can be either Gaussian-type or Bernoulli-type. A Gaussian-type unit can take continuous values ranging from minus infinity to plus infinity, whereas a Bernoulli-type unit is binary and takes only the values 0 or $1$. In this study, we adopt a Bernoulli-Bernoulli RBM, in which both the visible and hidden units are of the Bernoulli type (Yamashita et al., 2014).
@@
where, $\theta=\{w_{ij},b_{i},c_{i}\}$ are the model parameters (Zhang et al., 2018). The weights $\boldsymbol{W}=w_{ij}$ represents the interaction between the visible unit $v_{i}$ and the hidden unit $h_{j}$. The parameters $\boldsymbol{b}=b_{i}$ and $\boldsymbol{c}=c_{i}$ represent the biases of the visible and hidden layers, respectively. Given the model parameter $\theta$, the probability of the visible layer being in the state $\boldsymbol{v}$ is given by,
$p(\boldsymbol{v}|\theta)=\frac{\sum_{\boldsymbol{h}}\exp\{-E(\boldsymbol{v},\boldsymbol{h}|\theta)\}}{Z(\theta)},$ (2)
@@
$Z(\theta)=\sum_{\boldsymbol{v}}\sum_{\boldsymbol{h}}\exp\{-E(\boldsymbol{v},\boldsymbol{h}|\theta)\}.$ (3)
@@
In general Boltzmann machines, computing the gradient of this cost function is intractable, while in the case of RBMs, it can be efficiently carried out using the Contrastive Divergence (CD) method (Hinton, 2002). In this study, we adopt the CD method as the optimization technique. We implemented the RBM model using Python, and by utilizing CuPy, we were able to accelerate the computations through GPGPU processing (Nishino and Loomis, 2017). The RBM code developed in this study is available on GitHub (Kobayashi and Watanabe, 2025).
@@ 2.2 Dataset
For the training data, we adopted compositions by J. S. Bach. A total of 58 MIDI files were obtained from the Mutopia Project (The Mutopia Project, 2025), and each file was converted into a black-and-white image representation known as a piano roll. A piano roll is a two-dimensional representation of musical information, where the horizontal axis corresponds to time and the vertical axis corresponds to pitch. Notes are depicted as horizontal bars, with their positions and lengths indicating the timing and duration of each note, respectively. Each pixel value in the piano roll image is either 0 or 1, corresponding to the binary visible units of a Bernoulli-type RBM.
@@
The image size was fixed at $72\times 192$ pixels. The vertical dimension of 72 pixels corresponds to the pitch range from C1 to B6, where C1 denotes the C note in the first octave of a standard 88-key piano (i.e., the lowest C key), and B6 denotes the B note in the sixth octave, one semitone below the highest C (C8). The horizontal dimension of 192 pixels represents time, with 24 pixels corresponding to the duration of one quarter note. This resolution was chosen so that the horizontal pixel count would be divisible by 3, enabling the representation of triplet notes. Image size and the number of hidden units used for training are summarized in Table 1.
@@
Table 1: Image size and number of hidden units for training
Image size $72\times 192$
Number of hidden units 2048
@@ 2.3 Music Composition
6: By repeating steps 2 through 5 NN times, a binary vector is obtained in which exactly NN elements are set to 1.
@@ 3.3 Music Composition
An example of two-measure music generation using Algorithm 1 is shown in Fig. 6. The figure shows the visible states $\mathbf{v}_{t}$ at sampling steps $t=50,100,250,500,750,$ and $1000$. All images exhibit the structure of piano rolls. The time evolutions of the energy of the RBM during image generation is shown in Fig. 8. The energy decreases monotonically up to approximately 500 sampling steps, after which it begins to increase. This suggests that the RBM assigns higher energy when the number of active pixels (notes) is either too small or too large, implying the existence of an optimal number of notes that minimizes the energy. The energy reached its minimum at sampling step $t=557$. The corresponding piano roll is shown in Fig. 7. An analysis of this piano roll reveals that all notes appearing in the segment belong to the pitch-class set of B minor. In addition, the diatonic chord E minor, which is one of the diatonic triads in B minor, is present in the generated segment. The phrase also contains a stepwise motion C#-D-C#-B, which is musically natural in the context of the B-minor scale. These observations indicate that the piano roll generated by the RBM exhibits musically ordered structure in terms of overall pitch content, harmonic organization, and melodic motion.
An example of an eight-measure composition generated using Algorithm 2 is shown in Fig. 9. This image also exhibited a piano roll structure, similar to the two-measure images shown in Fig. 7. A close inspection of the piano roll shows that the pitch organization in measures 1-3 is based on the F-major key, exhibiting a musically ordered structure. However, after the third measure, the pitch content gradually becomes more irregular, and the musical coherence diminishes. Therefore, it is considered difficult for the trained RBM in its current form to generate piano rolls that exceed the number of measures in the training data while maintaining musically coherent structure. The audio of this piece, as well as other compositions generated by the RBM trained in this study, can be found online (rbm, ).
@@
(a)Refer to caption
(b)Refer to caption
Figure 7: A 2-measure piece composed by the RBM. (a) Piano roll representation of the 2-measure piece. (b) Sheet music of the 2-measure piece.
@@
Figure 9: An 8-measure piece composed by the RBM. (a) Piano roll representation of the 8-measure piece. (b) Sheet music of the 8-measure piece.
3.4 Analysis of Internal Representations
To investigate what kinds of patterns the RBM extracted from the musical training data, we provided one-hot vectors to the hidden layer of the trained model and computed the corresponding expected values of the visible layer. When these expected values were visualized as a colormap, numerous local temporal patterns with the width of a sixteenth-note duration were observed. This result indicates that the RBM spontaneously extracts elements corresponding to note duration from the input music. Therefore, the RBM can be regarded as having acquired internal representations that enable the reconstruction of rhythm structures based on note values.
On the other hand, typical melodic phrases or chordal structures were scarcely observed in the extracted patterns, suggesting that the internal representations of the trained RBM are not readily interpretable in terms of human musical intuition. It has been pointed out that the latent representations of standard RBMs often consist of complex mixtures of multiple features (Fernandez-de Cossio-Diaz et al., 2023), and it is therefore difficult for the model to acquire feature-separated internal representations—such as those corresponding to specific chordal or harmonic structures—without explicit label information.
Nevertheless, the fact that the RBM’s internal representations do not directly correspond to human music-theoretical concepts suggests that the model captures the statistical structure of musical data from a perspective fundamentally different from that of human music theory. In this sense, the latent space acquired by the RBM may serve as a data-driven analytical representation that does not rely on conventional theoretical frameworks. The results corresponding to this analysis are shown in Fig. 10.
(a)Refer to caption
(b)Refer to caption
[…]
(c)Refer to caption
Figure 10: Visualization of the expected visible-layer activations obtained by providing a one-hot vector that activates only the first hidden unit of the trained RBM. (a) Colormap representation of the expected activation pattern induced by this hidden unit. (b) The same colormap as in (a), with vertical grid lines added at sixteenth-note intervals to facilitate the identification of temporal structures. (c) Binary representation created by averaging the expected activations over each sixteenth-note interval and converting values of 0.55 or higher to white, with all lower values shown in black.
4 Summary and Discussion
We demonstrated that music composition is feasible even with a structurally simple model such as an RBM. By representing musical scores in piano-roll format, we enabled the model to learn musical features using techniques analogous to those employed in image modeling. The trained RBM successfully reconstructed piano-roll representations, including those derived from musical pieces not seen during training, while failing to reconstruct non-musical images and assigning high energy values to such inputs. Although the training data were limited to two-measure piano rolls, we developed a generation algorithm that allowed the model to produce musical sequences of arbitrary length.
The simplicity of the RBM architecture allowed us to analyze how the trained model internally represents musical data in a more direct manner than would be feasible with more complex models. By examining the hidden-layer activations in response to various inputs, we found that musical transposition caused substantial changes in the internal states, suggesting that the RBM evaluates musical similarity primarily based on the overlap of absolute pitch positions rather than on abstract melodic structure. This behavior is consistent with previous observations that RBMs and Deep Belief Networks lack inherent translational invariance in their input space. In contrast, convolutional deep belief models, which incorporate local receptive fields and weight sharing, offer a potential path toward improved recognition of transposed musical patterns due to their translational invariance (Lee et al., 2009).
Future work will extend the present framework to musical corpora beyond the works of J. S. Bach in order to examine whether training on different composers or musical genres leads to systematically distinct generative characteristics. Such studies may clarify whether restricted Boltzmann machines can extract and reproduce composer-specific or genre-specific stylistic features. In addition, recent studies have suggested that the tasks learned by RBMs are reflected in the singular value spectrum of their weight matrices (Ichikawa and Hukushima, 2022), and examining how the composer or genre influences this spectrum represents a promising direction for further research. Moreover, previous work has indicated that hidden units in RBMs can encode prototypical patterns in the visible layer (Hinton, 2002), although such patterns are often difficult to interpret in standard RBMs. Architectures such as classification RBMs, which tend to produce more distinguishable hidden-unit activations (Larochelle et al., 2012), may facilitate the identification of prototypical melodic or harmonic structures, and exploring such architectural extensions remains an important topic for future investigation.
Acknowledgment
This research was supported by JSPS KAKENHI, Grant No. JP21K11923. The computation was partly carried out using the facilities of the Supercomputer Center, Institute for Solid State Physics (ISSP), the University of Tokyo (ISSPkyodo-SC-2025-Cb-0038).
Appendix A Analysis of the Hidden Layer Using t-SNE
We confirmed that the RBM trained on piano-roll data is capable of music generation. To examine how the trained RBM internally represents musical structure, we analyzed its responses to transposed musical inputs. In an RBM, information presented to the visible layer is compressed and encoded in the hidden layer, from which the visible states can be reconstructed. The hidden layer was therefore examined to characterize internal representations of musical inputs.